# An LLL algorithm with symmetries

### v1

@@ 4. The case $\mathbf{G}=\mathrm{SO}_{2g}$
Let $g\geq 3$ be an integer, and let $e_{-g},e_{-1},e_{1},\dots,e_{g}$ denote the standard basis of $\mathbb{R}^{2g}$. Let $\Psi$ denote the $g\times g$ matrix with $1$’s on the antidiagonal and 00’s elsewhere, and let
@@
  • For each $i=-g+1,\dots,-1$, successively make the entries $n_{j,i}$ ($j=i-1,\dots,-g$) lie in $[-1/2,1/2)$ by multiplying on the right by $u_{\varepsilon_{i}-\varepsilon_{j}}(m_{i,j})$ for some integer $m_{i,j}$.
  • For each $i=-1,\dots,-g$, successively make the entries $n_{j,-i}$ ($j=i,\dots,-g$) lie in $[-1/2,1/2)$ by multiplying on the right by $u_{\varepsilon_{i}+\varepsilon_{j}}(m_{i,j})$ for some integer $m_{i,j}$.
@@
Algorithm 4.4.
Fix $\delta\in(1/4,1)$. This algorithm takes as input an inner product $H$ on $\mathbb{R}^{2g}$ compatible with $S$, and produces an element $\gamma\in\mathrm{SO}_{2g}(\mathbb{Z})$ such that ${}^{t}\gamma H\gamma$ corresponds to a $\delta$-reduced point of $X_{\mathrm{SO}_{2g}}$ under the bijection of Proposition 3.1.
  1. (1)
    (Initialise) Using Gram–Schmidt, compute $a\in A,n\in N$ such that $H={}^{t}(an)an$. Set $a^{\prime}\leftarrow a$, $n^{\prime}\leftarrow n$, $\gamma\leftarrow 1\in\mathrm{SO}_{2g}(\mathbb{Z})$, $k\leftarrow-g$.
  2. (2)
    (Iteration for $k<-1$) Execute RED($k,k+1$).
    1. (a)
      If $\alpha_{k}(a^{\prime})^{2}>(\delta-(n^{\prime}_{k,k+1})^{2})^{-1}$: execute REFL($k$). If $k>-g$, set $k\leftarrow k-1$ and go to Step 2. Otherwise, go to Step 2.
    2. (b)
      If $\alpha_{k}(a^{\prime})^{2}\leq(\delta-(n^{\prime}_{k,k+1})^{2})^{-1}$: execute RED($i,k+1$) for $i=k-1,\dots,-g$. Set $k\leftarrow k+1$. If $k<-1$, go to Step 2. If $k=-1$, go to Step 3.
  3. (3)
    (Iteration for $k=-1$) Execute RED($-2,1$).
    1. (a)
      If $\alpha_{k}(a^{\prime})^{2}>(\delta-(n^{\prime}_{-1,1})^{2})^{-1}$: execute REFL($-1$). Set $k\leftarrow k-2$ and go to Step 2.
    2. (b)
      If $\alpha_{k}(a^{\prime})^{2}\leq(\delta-(n^{\prime}_{-1,1})^{2})^{-1}$: execute RED($i,1$) for each $i=-3,\dots,-g$. For each $j=2,\dots,g$, execute RED($i,j$) for each $i=-j-1,\dots,-g$. Go to Step 4.
  4. (4)
    (End) Return $\gamma$.
  5. RED($i,j$)
    1. (a)
      If $-g\leq i<j\leq-1$: let $m$ be the unique integer such that $n^{\prime}_{i,j}+t\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{e_{i}-e_{j}}(m)$, $\gamma\leftarrow\gamma u_{e_{i}-e_{j}}(m)$.
    2. (b)
      If $-g\leq i<-j\leq-1$: let $m$ be the unique integer such that $n^{\prime}_{i,j}+t\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{e_{i}+e_{-j}}(m)$, $\gamma\leftarrow\gamma u_{e_{i}+e_{-j}}(m)$.
  6. REFL($i$)
    1. (a)
      If $-g\leq i\leq-2$: set $t\leftarrow n_{i,i+1}$, $m^{\prime}\leftarrow u_{e_{i}-e_{i+1}}(t)^{-1}n^{\prime}$, $n^{\prime}\leftarrow u_{e_{i}-e_{i+1}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})w_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.
    2. (b)
      If $i=-1$: set $t\leftarrow n_{-1,1}$, $m^{\prime}\leftarrow u_{2e_{i}}(t)^{-1}n^{\prime}$, $n^{\prime}\leftarrow u_{e_{-g+1}+e_{-g}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})w_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.
@@ 6. REFL($i$)
(More precisely, the proof of [SW18, Theorem 10] shows how rational points of the Jacobian of the curve given rise to $\mathbf{G}(\mathbb{Q})$-orbits; the proof of [SW18, Corollary 13] yields, in the simple case of orbits corresponding to integral points on the curve, an easy construction of an integral representative for this rational orbit.) The reduction covariant $H_{T}\in X_{\mathrm{SO}_{8}(\mathbb{R})}$ may be computed using the same method as in [Tho, §3]; we obtain the inner product on $\mathbb{R}^{8}$, compatible with $S$, given (to 6 significant figures) by the matrix
@@
Ideally we would like to generalise this example by allowing elements of $V(\mathbb{Z})$ corresponding to arbitrary rational points of the Jacobian of the hyperelliptic curve $Y^{2}=f(X)$. It is a relatively simple matter to write down elements of $V(\mathbb{Q})$ corresponding to such points (for example, by using a construction similar to the one given in [Tho14] for odd hyperelliptic curves), but showing these elements are $\mathbf{G}(\mathbb{Q})$-conjugate to points of $V(\mathbb{Z})$ requires a minimisation algorithm generalising those given in [CFS10] for elliptic curves. Such an algorithm has been given for curves of genus 2 by Fisher–Liu [FY23], although in this case our algorithm is not needed due to the existence of an exceptional isomorphism $\mathrm{PSO}_{6}\cong\mathrm{PGL}_{4}$ (see [FL23, Remark 4.3]), a phenomenon that does not persist when $62g>$ 6.
@@ 5. The case $\mathbf{G}=G_{2}$
$u_{\alpha_{1}}(t)=\left(\begin{array}[]{cccccccc}1&t&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&t&-t&t^{2}&0&0\\ 0&0&0&1&0&t&0&0\\ 0&0&0&0&1&-t&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&-t\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),u_{\alpha_{2}}(t)=\left(\begin{array}[]{cccccccc}1&0&0&0&0&0&0&0\\ 0&1&t&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&-t&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),u_{\alpha_{3}}(t)=\left(\begin{array}[]{cccccccc}1&0&t&0&0&0&0&0\\ 0&1&0&-t&t&0&t^{2}&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&-t&0\\ 0&0&0&0&1&0&t&0\\ 0&0&0&0&0&1&0&-t\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),$
$u_{\alpha_{4}}(t)=\left(\begin{array}[]{cccccccc}1&0&0&t&-t&0&0&t^{2}\\ 0&1&0&0&0&-t&0&0\\ 0&0&1&0&0&0&t&0\\ 0&0&0&1&0&0&0&t\\ 0&0&0&0&1&0&0&-t\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),u_{\alpha_{5}}(t)=\left(\begin{array}[]{cccccccc}1&0&0&0&0&t&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&-t\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),u_{\alpha_{6}}(t)=\left(\begin{array}[]{cccccccc}1&0&0&0&0&0&t&0\\ 0&1&0&0&0&0&0&-t\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right).$
@@
$s_{1}=\left(\begin{array}[]{cccccccc}0&1&0&0&0&0&0&0\\ -1&0&0&0&0&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&0&0&0&0&-1\\ 0&0&0&0&0&0&1&0\\ \end{array}\right),s_{2}=\left(\begin{array}[]{cccccccc}1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&-1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&-1&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right).$
@@
Algorithm 5.4.
Fix $\delta\in(1/4,1)$. This algorithm takes as input an inner product $H$ on $\mathbb{O}_{\mathbb{R}}$ compatible with $q$ and $\cdot$, and produces an element $\gamma\in G_{2}(\mathbb{Z})$ such that ${}^{t}\gamma H\gamma$ corresponds to a $\delta$-reduced point of $X_{G_{2}}$ under the bijection of Proposition 5.1.
  1. (1)
    (Initialise) Using Gram–Schmidt, compute $a\in A,n\in N$ such that $H={}^{t}(an)an$. Set $a^{\prime}\leftarrow a$, $n^{\prime}\leftarrow n$, $\gamma\leftarrow 1\in G_{2}(\mathbb{Z})$.
  2. (2)
    Execute RED($1$).
    1. (a)
      If $\alpha_{1}(a^{\prime})^{2}>(\delta-(n^{\prime}_{1,2})^{2})^{-1}$: execute REFL($1$) and go to Step 2.
    2. (b)
      If $\alpha_{1}(a^{\prime})^{2}\leq(\delta-(n^{\prime}_{1,2})^{2})^{-1}$: go to Step 3.
  3. (3)
    Execute RED($2$).
    1. (a)
      If $\alpha_{2}(a^{\prime})^{2}>(\delta-(n^{\prime}_{2,3})^{2})^{-1}$: execute REFL($2$) and go to Step 2.
    2. (b)
      If $\alpha_{2}(a^{\prime})^{2}\leq(\delta-(n^{\prime}_{2,3})^{2})^{-1}$: execute RED($i$) for each $i=3,\dots,6$ and then go to Step 4.
  4. (4)
    (End) Return $\gamma$.
  5. RED($i$)
    Let $(k,l)$ be the $i^{\text{th}}$ tuple in the list
    $(1,2),(2,3),(1,3),(1,4),(1,6),(1,7),$
    and let $t$ be the unique integer such that $n^{\prime}_{k,l}+t\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow n^{\prime}u_{\alpha_{i}}(-t)$, $\gamma\leftarrow\gamma u_{\alpha_{i}}(t)$.
  6. REFL($i$)
    Let $(k,l)$ be the $i^{\text{th}}$ tuple in the list $(1,2),(2,3)$. set $t\leftarrow n_{k,l}$, $m^{\prime}\leftarrow u_{\alpha_{i}}(-t)n^{\prime}$, $n^{\prime}\leftarrow u_{\alpha_{i}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})w_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.

### v2

@@ 4. The case $\mathbf{G}=\mathrm{SO}_{2g}$
Let $g\geq 3$ be an integer, and let $e_{-g},\dots,e_{-1},e_{1},\dots,e_{g}$ denote the standard basis of $\mathbb{R}^{2g}$. Let $\Psi$ denote the $g\times g$ matrix with $1$’s on the antidiagonal and 00’s elsewhere, and let
@@
Given an inner product $H$ compatible with $S$, let $n=n_{H}=(\mu_{i,j})_{-g\leq i,j\leq g,ij\neq 0}$ and $a=a_{H}=\operatorname{diag}(\|e_{i}^{\ast}\|_{H})_{-g\leq i\leq g,i\neq 0}$ as defined above. Then $a\in A$, $n\in N$, and $H={}^{t}(an)an$.
@@
  • For each $j=-g+1,\dots,-2,-1$, successively make the entries $n_{i,j}$ (for $i=j-1,j-2,\dots,-g$) lie in $[-1/2,1/2)$ by multiplying on the right by $u_{\varepsilon_{i}-\varepsilon_{j}}(m_{i,j})$ for some integer $m_{i,j}$.
  • For each $j=1,2,\dots,g$, successively make the entries $n_{i,j}$ (for $i=-j-1,-j-2,\dots,-g$) lie in $[-1/2,1/2)$ by multiplying on the right by $u_{\varepsilon_{i}+\varepsilon_{-j}}(m_{i,j})$ for some integer $m_{i,j}$.
@@
Algorithm 4.4 Lattice reduction for $\mathrm{SO}_{2g}$.
Input: A constant $\delta\in(1/4,1)$, and a positive definite $2g\times 2g$ real symmetric matrix $H$ compatible with $S$.
Output: An element $\gamma\in\mathrm{SO}_{2g}(\mathbb{Z})$ such that ${}^{t}\gamma H\gamma$ corresponds to a $(\delta,\omega)$-reduced point of $X_{\mathrm{SO}_{2g}(\mathbb{R})}$ under the bijection of Proposition 4.1.
1:  Using Gram–Schmidt, compute $a\in A,n\in N$ such that $H={}^{t}(an)an$. Set $\gamma\leftarrow 1\in\mathrm{SO}_{2g}(\mathbb{Z})$, $a^{\prime}\leftarrow a$, $n^{\prime}\leftarrow n$, $k\leftarrow-g$.
2:  while $k<-1$ do
3:     Execute RED($k,k+1$).
4:     if  $\alpha_{k}(a^{\prime})^{2}>(\delta-(n^{\prime}_{k,k+1})^{2})^{-1}$  then
5:        Execute REFL($k$).
6:        if $k>-g$ then
7:           $k\leftarrow k-1$.
8:        end if
9:     else
10:        for  $i=k-1$, $i\geq-g$, $i\leftarrow i-1$  do
11:           Execute RED($i,k+1$).
12:        end for
13:        Set $k\leftarrow k+1$.
14:     end if
15:  end while
16:  Execute RED($-2,1$).
17:  if $\alpha_{-1}(a^{\prime})^{2}>(\delta-(n^{\prime}_{-1,1})^{2})^{-1}$ then
18:     Execute REFL($-1$). Set $k\leftarrow k-2$.
19:     go to 2.
20:  else
21:     for $i=-3$, $i\geq-g$, $i\leftarrow i-1$ do
22:        Execute RED($i,1$).
23:     end for
24:     for $j=2$, $j\leq g$, $j\leftarrow j+1$ do
25:        for $i=-j-1$, $i\geq-g$, $i\leftarrow i-1$ do
26:           Execute RED($i,j$).
27:        end for
28:     end for
29:  end if
30:  return  $\gamma$.
Algorithm 4.5 Subroutine RED($i,j$) for Algorithm 4.4.
  if  $-g\leq i<j\leq-1$  then
     Let $m$ be the unique integer such that $n^{\prime}_{i,j}+m\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{\varepsilon_{i}-\varepsilon_{j}}(m)$, $\gamma\leftarrow\gamma u_{\varepsilon_{i}-\varepsilon_{j}}(m)$.
  else if  $-g\leq i<-j\leq-1$ then
     Let $m$ be the unique integer such that $n^{\prime}_{i,j}+m\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{\varepsilon_{i}+\varepsilon_{-j}}(m)$, $\gamma\leftarrow\gamma u_{\varepsilon_{i}+\varepsilon_{-j}}(m)$.
  end if
Algorithm 4.6 Subroutine REFL($i$) for Algorithm 4.4.
  if $-g\leq i\leq-2$ then
     Set $t\leftarrow n_{i,i+1}$, $m^{\prime}\leftarrow u_{\varepsilon_{i}-\varepsilon_{i+1}}(t)^{-1}n^{\prime}$, $n^{\prime}\leftarrow u_{\varepsilon_{i}-\varepsilon_{i+1}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})s_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.
  else if $i=-1$ then
     Set $t\leftarrow n_{-1,1}$, $m^{\prime}\leftarrow u_{2\varepsilon_{i}}(t)^{-1}n^{\prime}$, $n^{\prime}\leftarrow u_{\varepsilon_{-g+1}+\varepsilon_{-g}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})s_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.
  end if
@@
(More precisely, the proof of [SW18, Theorem 10] shows how rational points of the Jacobian of the curve give rise to $\mathbf{G}(\mathbb{Q})$-orbits; the proof of [SW18, Corollary 13] yields, in the simple case of an orbit corresponding to an integral point on the curve, an easy construction of a rational representative for this orbit whose denominators are bounded powers of $2$.) The reduction covariant $H_{T}\in X_{\mathrm{SO}_{8}(\mathbb{R})}$ may be computed using the same method as in [Tho, §3]; we obtain the inner product on $\mathbb{R}^{8}$, compatible with $S$, given (to 6 significant figures) by the matrix
@@
Ideally we would like to generalise this example by allowing elements of $V(\mathbb{Z})$ corresponding to arbitrary rational points of the Jacobian of the hyperelliptic curve $Y^{2}=f(X)$. It is a relatively simple matter to write down elements of $V(\mathbb{Q})$ corresponding to such points (for example, by using a construction similar to the one given in [Tho14] for odd hyperelliptic curves), but showing these elements are $\mathbf{G}(\mathbb{Q})$-conjugate to points of $V(\mathbb{Z})$ requires a minimisation algorithm generalising those given in [CFS10] for elliptic curves. Such an algorithm has been given for curves of genus 2 by Fisher–Liu [FL23], although in this case our algorithm is not needed due to the existence of an exceptional isomorphism $\mathrm{PSO}_{6}\cong\mathrm{PGL}_{4}$ (see [FY23, Remark 4.3]), a phenomenon that does not persist when $62g>$ 6.
@@ 5. The case $\mathbf{G}=G_{2}$
$u_{\alpha_{1}}(t)=\left(\begin{array}[]{cccccccc}1&t&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&t&-t&t^{2}&0&0\\ 0&0&0&1&0&t&0&0\\ 0&0&0&0&1&-t&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&-t\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),\enspace u_{\alpha_{2}}(t)=\left(\begin{array}[]{cccccccc}1&0&0&0&0&0&0&0\\ 0&1&t&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&-t&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),\enspace u_{\alpha_{3}}(t)=\left(\begin{array}[]{cccccccc}1&0&t&0&0&0&0&0\\ 0&1&0&-t&t&0&t^{2}&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&-t&0\\ 0&0&0&0&1&0&t&0\\ 0&0&0&0&0&1&0&-t\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),$
$u_{\alpha_{4}}(t)=\left(\begin{array}[]{cccccccc}1&0&0&t&-t&0&0&t^{2}\\ 0&1&0&0&0&-t&0&0\\ 0&0&1&0&0&0&t&0\\ 0&0&0&1&0&0&0&t\\ 0&0&0&0&1&0&0&-t\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),\enspace u_{\alpha_{5}}(t)=\left(\begin{array}[]{cccccccc}1&0&0&0&0&t&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&-t\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),\enspace u_{\alpha_{6}}(t)=\left(\begin{array}[]{cccccccc}1&0&0&0&0&0&t&0\\ 0&1&0&0&0&0&0&-t\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right).$
@@
$s_{1}=\left(\begin{array}[]{cccccccc}0&1&0&0&0&0&0&0\\ -1&0&0&0&0&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&0&0&0&0&-1\\ 0&0&0&0&0&0&1&0\\ \end{array}\right),\quad s_{2}=\left(\begin{array}[]{cccccccc}1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&-1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&-1&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right).$
@@
Algorithm 5.4 Lattice reduction for $G_{2}$.
Input: A constant $\delta\in(1/4,1)$, and an inner product $H$ on $\mathbb{O}_{\mathbb{R}}$ compatible with $q$ and $\cdot$.
Output: An element $\gamma\in G_{2}(\mathbb{Z})$ such that ${}^{t}\gamma H\gamma$ corresponds to a $\delta$-reduced point of $X_{G_{2}}$ under the bijection of Proposition 5.1.
1:  Using Gram–Schmidt, compute $a\in A,n\in N$ such that $H={}^{t}(an)an$. Set $a^{\prime}\leftarrow a$, $n^{\prime}\leftarrow n$, $\gamma\leftarrow 1\in G_{2}(\mathbb{Z})$.
2:  Execute RED($1$).
3:  if $\alpha_{1}(a^{\prime})^{2}>(\delta-(n^{\prime}_{1,2})^{2})^{-1}$ then
4:     Execute REFL($1$).
5:     go to 2.
6:  end if
7:  Execute RED($2$).
8:  if $\alpha_{2}(a^{\prime})^{2}>(\delta-(n^{\prime}_{2,3})^{2})^{-1}$ then
9:     Execute REFL($2$).
10:     go to 2.
11:  else
12:     for $i=3$ to $6$ do
13:        Execute RED($i$).
14:     end for
15:  end if
16:  return  $\gamma$.
Algorithm 5.5 Subroutine RED($i,j$) for Algorithm 5.4.
  Let $(k,l)$ be the $i^{\text{th}}$ tuple in the list $(1,2),(2,3),(1,3),(1,4),(1,6),(1,7),$ and let $m$ be the unique integer such that $n^{\prime}_{k,l}+m\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow n^{\prime}u_{\alpha_{i}}(-m)$, $\gamma\leftarrow\gamma u_{\alpha_{i}}(m)$.
Algorithm 5.6 Subroutine REFL($i$) for Algorithm 5.4.
  Let $(k,l)$ be the $i^{\text{th}}$ tuple in the list $(1,2),(2,3)$. Set $t\leftarrow n_{k,l}$, $m^{\prime}\leftarrow u_{\alpha_{i}}(-t)n^{\prime}$, $n^{\prime}\leftarrow u_{\alpha_{i}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})s_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.